OpenStudy (anonymous):

I NEED AN ANSWER WITH WORK SHOWN ⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵⤵✔ Evaluate the square root of forty-nine plus four squared minus three minus the square root of twenty-five. •-5 • 3 •15 • 25

OpenStudy (solomonzelman):

allcharacters.com or whatever the name of the website is (?)

OpenStudy (anonymous):

im on florida virtual 8th grade math

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ \sqrt{49}+4^2-\sqrt{25}}\) I would think this is the question.

OpenStudy (solomonzelman):

The site I was talking about, I meant that you got your symbols from there... anyway, back to the problem. \(\normalsize\color{blue}{ \sqrt{49}+4^2-\sqrt{25}}\) you take the potivie values of the roots, \(\normalsize\color{blue}{ 7+4^2-5}\) and 4²=16, so \(\normalsize\color{blue}{ 7+16-5}\)

OpenStudy (solomonzelman):

your answer is ?

OpenStudy (solomonzelman):

Who closed the question, armistre, or you ?